4. Three times the square of a certain positive number exceeds six times the number by nine. Find the number:
a. 3 b. 6 c. 4 d. 9 e. none
The answer is A, but how? I don't think I used the correct formula. Please help!

Answer by fractalier(6550)

Let the number be x. Then we have
3x^2 - 6x = 9
3x^2 - 6x - 9 = 0
x^2 - 2x - 3 = 0
(x - 3)(x + 1) = 0
x = 3 and x = -1
but since you need positive only...
x = 3
Choice A.